The package easytable

Enrico Bertolazzi

Laboratorio di Matematica Applicata e Meccanica Computazionale

Dipartimento di Ingegneria Meccanica e Strutturale

Università degli Studi di Trento Mesiano, Trento, Italy

enrico.bertolazzi@ing.unitn.it

22nd October 1999

Abstract

The easytable package is a macro package for writing tables, with equal column widths or equal rows heights or both, with various kinds of rules (lines) between rows and columns. It uses an array/tabular-like syntax.

1 How to use it

Load the package in the usual way:

\documentclass{article}
.
.
\usepackage[thinlines,thiklines]{easytab}
.
.

The options thinlines, and thiklines are self explanatory.